OpenStudy (anonymous):

evaluate log base 9 of (1/27) do u divide 9 into the fraction?

OpenStudy (jhannybean):

\[\log_9 \left(\frac{1}{27}\right) = x\] Now you're looking for a value of x that will turn 9 into 1/27, so \[\ 9^x=\frac{1}{27}\]

OpenStudy (anonymous):

it would be a fraction?

OpenStudy (jhannybean):

Remember, if x is a number, then \(\ x^{-1} = \frac{1}{x}\)

OpenStudy (anonymous):

so it would be -3?

OpenStudy (zarkon):

\[\log_{9}(1/27)=-\log_{9}(27)=-\frac{\log_3(27)}{\log_{3}(9)} \]

OpenStudy (jhannybean):

\[\ 9^{-3} = \frac{1}{729}\]

OpenStudy (anonymous):

so it would be a fraction?

OpenStudy (jhannybean):

But look! I just realized something, haha. 9 is a factor of 729, is it not?

OpenStudy (jdoe0001):

\(\bf \large { log_{\color{red}{ a}}{\color{blue}{ b}}=y\iff {\color{red}{ a}}^y={\color{blue}{ b}}\qquad thus \\ \quad \\ log_{\color{red}{ 9}}{\color{blue}{\frac{1}{27} }}=\square\implies {\color{red}{ 9}}^\square ={\color{blue}{ \frac{1}{27}}}\quad thus \quad \square =? \\ \quad \\ \textit{bear in mind that } \cfrac{1}{a^{\color{red} n}}\implies a^{-{\color{red} n}}\quad thus \\ \quad \\ 9^\square =\cfrac{1}{27} \implies (3^2)^\square =\cfrac{1}{3^3}\implies 3^{2\square }=3^{-3} }\)

OpenStudy (anonymous):

it is so what if it is though, from what this guy is saying, it would be the base with a exponent

OpenStudy (anonymous):

i did this in wolframalpha, and it came up with -3/2

OpenStudy (jhannybean):

I got it. I figured out an alternative way of explaining how @Zarkon derived the function.

OpenStudy (anonymous):

ok

OpenStudy (jhannybean):

Using logarthmic rules, we've concluded \[9^x = \frac{1}{27}\]3 is a common factor between both of these, and we know in order to evaluate functions that are powers, we need to same base. rewrite 1/27 as 27^{-1} \[9^x = 27^{-1}\] Now set the base to 3. 9 3^2 and 27 is 3^3. \[ (3^2)^x = (3^3)^{-1}\]\[ 3^{2 \cdot x} = 3^{-1 \cdot 3}\]\[3^{2x} = 3^{-3}\]\[\color{red}3^{2x} = \color{red}3^{-3} \therefore \]\[2x = -3\] Can you solve for x now? :)

OpenStudy (jhannybean):

Sorry, typo. 9 = 3^2*

OpenStudy (anonymous):

oh ok so thts how it equals -3/2

OpenStudy (anonymous):

ok, thanks

OpenStudy (jhannybean):

Yeah, I don't quite remember how to solve using lug rules so I used what knowledge I had of log functions. took a little while to remember though!

OpenStudy (anonymous):

well let me just say you're doing a fabulous job cause i just learned this today and i have a quiz on it tomorrow

OpenStudy (anonymous):

so using this, log base 8 of 32 would be 4 right?

OpenStudy (jhannybean):

Practice these problems you're posting up on your own without looking at the solutions people have posted up here. That way it'll be ingrained into your head for tomorrow :)